OpenStudy (a1234):

How do I write sin(x+y)cos(x-y) as a sum or difference using the product-to-sum formulas?

OpenStudy (solomonzelman):

the rules are: \(\large\color{black}{ \displaystyle \sin(a\pm b)=\sin(a)\cos(b)\pm\sin(b)\cos(a)}\) \(\large\color{black}{ \displaystyle \cos(a\pm b)=\cos(a)\cos(b)\mp\sin(a)\sin(b)}\)

OpenStudy (a1234):

I did those and came up with (sinx cosy + cosx siny) (cosx cosy + sinx siny)

OpenStudy (solomonzelman):

Ok, very good!

OpenStudy (a1234):

But the choices I have are: A. 1/2(sin2x + sin2y) B. -1/2(sin2x + sin2y) C. 2sinxsiny D. -2sinxsiny

OpenStudy (a1234):

@SolomonZelman

OpenStudy (solomonzelman):

expand it and try to re-write it. What you had is correct, but in different form.